# Amortized Sampling with Transferable Normalizing Flows

### v2

@@ 1 Introduction
Deep learning-based samplers, although considered in different settings, abandon the Markov chain approach to generating samples and shift the computational burden to a one-time training phase, enabling fast and inexpensive inference compared to MCMC. In the most challenging scenario, these methods consider having access only to the unnormalized density function (analogous to MC methods) (Vargas et al., 2023; Akhound-Sadegh et al., 2024). Boltzmann Generators (BGs) (Noé et al., 2019) consider a more practical scenario when, in addition to the unnormalized density, a dataset of MD trajectories is available, which does not necessarily match the target density. To mitigate the error introduced by the imperfections of the model and training data, BGs rely on training likelihood-based models and perform self-normalized importance sampling (SNIS) (Liu, 2001) at inference time. The availability of the training data and inference-time IS have enabled BGs to generalize across dipeptide systems (Klein and Noe, 2024), but they have not yet been able to generalize across larger and more diverse systems of scientific interest.
In this work, we introduce Prose, a large-scale normalizing flow which demonstrates unprecedented ability to transfer to previously unseen systems of varying amino acid composition, sequence length, and temperatures, outperforming MD for the same computational budget (see Fig.˜1). Our approach is strikingly simple and scalable, which elucidates the potential of the deep learning-based samplers for sampling applications. In particular, we achieve this through the following series of contributions:
  • We introduce ManyPeptidesMD: a novel dataset of molecular dynamics trajectories for peptide systems between 2 and 8 residues. The training dataset consists of 21,700 peptide sequences simulated for 200 ns each, with 10,000 sequences for octopeptides alone.
@@ 2.1 Normalizing flows
The fundamental challenge of probabilistic modeling is designing a density model from which samples can be efficiently generated. Normalizing flows (Rezende and Mohamed, 2015) approach this challenge by defining a diffeomorphism — differentiable invertible function with a differentiable inverse. Namely, given a simple prior density $q_{z}(z)$ and a parameterized flow (diffeomorphism) $f_{\theta}(x)$, one can define the push-forward distribution as the map of samples from the prior distribution $z\sim q_{z}(z)$ via the inverse flow $x=f^{-1}_{\theta}(z)\sim q_{\theta}(x)$ with learnable parameters $\theta$. The density of the push-forward distribution can then be computed via the change-of-variables formula
@@
Auto-regressive normalizing flows (Kingma et al., 2016; Papamakarios et al., 2017; Zhai et al., 2024) define a family of invertible maps with tractable Jacobian as a sequence of composed transformations $f_{\theta}=f_{\tau}\circ\ldots\circ f_{0}$, where each transformation $z_{t+1}=f_{t}(z_{t})$ is defined auto-regressively. In the case of the TarFlow (Zhai et al., 2024), this auto-regression is an affine update defined over image patches $x[i]\in\mathbb{R}^{C\times H\times W}$ and parameterized by transformer blocks. That is, the $i$-th token of the output is
@@
where we adopt slicing notation denoting the $i$-th dimension as $z[i]$ and all the dimensions up to $i$-th (exclusive) as $z[:i]$. Notably, the auto-regressive structure allows for efficient evaluation of the Jacobian determinant due to its lower-triangular structure $\log|\partial f_{t}(z_{t})/\partial{z_{t}}|=-\sum_{i=1}^{d-1}\alpha_{t}(z_{t}[:i])[i]$. Furthermore, the auto-regressive affine updates are invertible with inverse $z_{t}=f^{-1}_{t}(z_{t+1})$ given by
$z_{t}[i]=\begin{cases}z_{t+1}[i]\,,&i=0\,,\\ z_{t+1}[i]\cdot\exp(\alpha_{t}(z_{t}[:i])[i])+\mu_{t}(z_{t}[:i])[i]\,,&i\in[1,d-1]\,.\end{cases}$ (3)
However, clearly, such transformations leave the leading dimension $z_{t}[0]$ untouched, hence must be interleaved with permutations over the dimensions $f_{\theta}=\pi_{\tau}\circ f_{\tau}\circ\ldots\circ\pi_{0}\circ f_{0}$, where $\pi_{t}$ is the permutation. For example, Zhai et al. (2024) use simple inversions for all $\pi_{t}$ across the entire model.
Normalizing flows are versatile probabilistic models, which allow various training strategies, such as variational inference for learning unnormalized target densities (Rezende and Mohamed, 2015), or maximum likelihood estimation for generative modeling (Kingma and Dhariwal, 2018). Analogously, one can use the model density $q_{\theta}(x)$ at inference time, which we discuss in the following section.
@@ 2.2 Boltzmann generators
Despite normalizing flows allowing various forms of supervision during training (e.g. unnormalized densities or empirical distributions), errors present in the generated samples prevents accurate evaluation within scientific applications requiring high precision, e.g. free energy estimation. Boltzmann generators address specifically this challenge by performing self-normalized importance sampling (SNIS) at inference time. Namely, to evaluate the expectation of a statistic $\varphi(x)$ w.r.t. the target Boltzmann density $p(x)$ one can use the following consistent Monte Carlo estimator
@@
where $q_{\theta}(x)$ is the density of the learned normalizing flow. The SNIS estimator converges, for $n\to\infty$, to the true value $\mathbb{E}_{p(x)}\varphi(x)$. Furthermore, Tan et al. (2025) demonstrated that one can reduce the variance of the SNIS estimator even further by leveraging the scalability of the recently proposed TarFlow (Zhai et al., 2024) and combining it with the continuous-time formulation of Sequential Monte Carlo (Jarzynski, 1997; Albergo and Vanden-Eijnden, 2025).
Transferable Boltzmann generator (TBG) (Klein and Noe, 2024) made a first attempt of learning a sampler that generalizes across target densities corresponding to different peptide systems. TBG parameterizes the proposal distribution as a continuous normalizing flow (CNF) (Chen et al., 2018b), where the vector field is defined by an equivariant graph neural network (Satorras et al., 2021; Klein et al., 2023c). The crucial part of the method is the system-dependent embedding of $N$ atoms
$h[i]=[x[i],A_{i},R_{i},P_{i}]\,$, (5)
where $h[i]$ corresponds to $i$-th atom in the system with the spatial coordinates $x[i]\in\mathbb{R}^{3}$ and system conditioning information $A_{i},R_{i},P_{i}]$ encoded as one-hot vectors: atom type $A_{i}$, residue type $R_{i}$, and position $P_{i}$. Training on a set of MD trajectories for dipeptide systems (Klein et al., 2023b) with this system-conditional encoding enables TBG to generate a proposal for previously unseen dipeptides.
Despite successful generalization across dipeptides, the TBG architecture introduces significant bottlenecks for inference and fine-tuning. Indeed, the learned CNF requires accurate integration of the vector field and computationally expensive evaluation of its divergence for evaluating the learned density model. For instance, the implementation of (Klein and Noe, 2024) requires 4 GPU-days to produce $3\times 10^{4}$ samples with their corresponding proposal likelihoods for a single dipeptide system. Furthermore, the expensive evaluation of density makes it infeasible to train or finetune TBG via the reverse KL-divergence or create a replay buffer of a substantial size.
@@ 3.1 Architecture of Prose
Prose builds on the recent TarFlow architecture (Zhai et al., 2024), which parameterizes a sequence of autoregressive affine transformations via blocks of transformer layers. The expressivity and favorable scalability of the transformer layers enables TarFlow to effectively model high dimensional data, whilst the affine autoregressive flow parameterization ensure fast and accurate energy evaluation. With minimal modifications TarFlow is capable of successfully modeling high-dimensional molecular data (Tan et al., 2025). Here we describe our design choices that make transferability possible.
@@
We extend TarFlow to support concurrent training on sequences of variable length. Whilst transformers natively support sequences of arbitrary length, special consideration is required within a normalizing flow such as TarFlow that is defined for fixed input and output dimensions. We therefore define appropriate masking to the affine sequence updates and log-determinant aggregation to prevent padding tokens influencing either computation, under arbitrary sequence permutations; further details of the tokenization and masking are provided in Appendix˜B. We additionally replace the fixed-length learnable position embedding with the more extrapolation-friendly sinusoidal embedding. This design enables Prose to efficiently train across a distribution of systems $s$ by maximizing the normalized log-likelihood
@@
where $d(s)$ is the size of the system $s$. This extended architecture allows for parallel processing of data dimensions, enabling transferability and scalability across lengths.
@@
The standard TarFlow employs simple additive conditioning for class-conditional image generation. Whilst we find this to be sufficient to define a system-transferable normalizing flow, we follow many large-scale atomistic transformer architectures in applying conditioning through adaptive layer normalization, adaptive scaling, and SwiGLU-based transition blocks (Abramson et al., 2024; Geffner et al., 2024). The system conditioning features are constructed from atom types $A$, residue types $R$, residue positions $P$, and sequence lengths $L$. Atom and residue types are embedded using lookup-table embedding layers, whilst sinusoidal embeddings are employed for the naturally ordered sequence position and sequence length. See Appendix˜B for further details.
@@
In the image setting, Zhai et al. (2024) employ only an identity and flip permutation to the sequence of image tokens. Similarly, when applying TarFlow to peptide systems Tan et al. (2025) employ only an identity and flip permutation on the ordering defined by the PDB format, in which the sequence is defined per-residue starting with backbone atoms followed by sidechain atoms. Whilst a simple identity and flip may be appropriate for the regular grid of image data, we argue this to be suboptimal for the diversity of geometric interactions present in molecular systems. This motivates our introduction of chemistry-aware sequence permutations, defined to promote effective peptide modeling. We define the backbone permutation, such that the backbone atoms $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=1}^{L}$ (with associated hydrogens) for all residues are located at the start of the sequence, and followed by the sidechains. By processing the coordinates of the backbone atoms at the start of the sequence, the model refines the global structure of the peptide as a contiguous sequence. Crucially, the sidechain positions are subsequently able to causally attend to the full backbone structure, hence enabling local updates to be influenced by global structure. We further employ a backbone-flip permutation to provide additional diversity to the autoregressive modeling.
Figure 2: All-atom block-wise autoregressive normalizing flow based on the TarFlow (Zhai et al., 2024). Peptides are encoded via atom types $A$, residue types $R$, residue position $P$, and sequence length $L$. Atom positions in 3D Cartesian coordinates define the system state. The embedding of the peptide is applied as conditioning to the coordinates such that Prose achieves transferability between systems. Within each block the sequence $z_{t}$ is permuted and passed to a transformer, defining an autoregressive affine update. In the backbone permutation the backbone $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=1}^{L}$ of all residues (with associated hydrogens) is updated before any sidechains, providing additional diversity to the causal attention for global structure modeling.
@@ 3.2 Inference and fine-tuning of Prose
In practice, we compare the true density $p(x)$ with our generated distribution $\tilde{p}(x)$ instead of measuring statistics $\varphi(x)$. For completeness, we describe all the considered estimators in Appendix˜C.
@@ 4.2 Experimental configuration
We train the first Boltzmann generators transferable across peptide sequence length. We train the Prose architecture defined in Section˜3.1, an unmodified TarFlow (Zhai et al., 2024) as in SBG (Tan et al., 2025), and the equivariant continuous normalizing flow of Klein and Noe (2024), with the improved training recipe of Tan et al. (2025), denoted as ECNF++. All models are trained for $5\times 10^{5}$ iterations with batch size 512. Both Prose and TarFlow are suitably scalable to long sequences and are trained on the full dataset detailed in Section˜4.1. However, generating 8 residue sequences with likelihoods for ECNF++ was found to be prohibitively expensive, hence the training data was limited to sequences up to and including length 4. Comprehensive training details are provided in Appendix˜D, results for the unweighted proposal distributions are provided in Appendix˜F.
@@ 4.3 Scale transferability of Prose
To establish the performance of Prose as a sampler proposal distribution, we first evaluate the trained flows in the Boltzmann generator setting. Here we generate a set of proposal particles $\{x_{i}\}_{i=1}^{N}$, evaluate model likelihoods $q_{\theta}(x_{i})$ and reweight using SNIS as in Eq.˜7. In addition to the trained models we benchmark against the following pretrained baselines; (i) the TBG model trained by Klein and Noe (2024), denoted as ECNF (ii) TimeWarp (Klein et al., 2023a) (iii) BioEmu (Lewis et al., 2024) (iv) Unisim (Yu et al., 2025). For all methods we permit a budget of $10^{4}$ energy evaluations. For the Boltzmann generator methods (ECNF, ECNF++, TarFlow, Prose) this corresponds to $10^{4}$ SNIS particles; for further information on the budget allocation of non-BG methods (TimeWarp, BioEmu, UniSim) see Section˜E.4. We note the TimeWarp dataset to lack any sequences containing Proline at the N-Terminal, hence neither TimeWarp nor ECNF were evaluated on such sequences.

### v3

@@ 1 Introduction
Deep learning-based samplers abandon the Markov chain approach to drawing samples and shift the computational burden to a one-time training phase, enabling fast and inexpensive inference compared to MCMC. In the most challenging scenario, these methods consider having access only to the unnormalized density function (analogous to MC methods) (Vargas et al., 2023; Akhound-Sadegh et al., 2024). Boltzmann generators (BGs) (Noé et al., 2019) consider a more practical scenario where, in addition to the unnormalized density, a dataset of MD trajectory is available, which does not necessarily match the target density. To mitigate the error introduced by imperfections of the model and training data, BGs train likelihood-based models and perform self-normalized importance sampling (SNIS) (Liu, 2001) at inference time. The availability of training data coupled with SNIS have enabled BGs to generalize across dipeptide systems (Klein and Noe, 2024), but they have not yet been able to generalize across larger and more diverse systems of scientific interest.
In this work, we introduce Prose, a large-scale normalizing flow which demonstrates unprecedented ability to transfer to previously unseen systems of varying amino acid composition, sequence length, and temperatures, outperforming MD for the same computational budget (see Fig.˜1). Our approach is strikingly simple and scalable, which elucidates the potential of the deep learning-based samplers for sampling applications. In particular, we outline the following series of contributions:
  • We introduce ManyPeptidesMD: a novel dataset of molecular dynamics trajectories for peptide systems between 2 and 8 residues. The training dataset consists of 21,700 peptide sequences simulated for $200\text{\,}\mathrm{ns}$ each, giving a total of $4.3\text{\,}\mathrm{ms}$ of simulation.
@@ 2.1 Normalizing flows
The fundamental challenge of probabilistic modeling is designing a density model from which samples can be efficiently generated. Normalizing flows (Rezende and Mohamed, 2015) approach this challenge by defining a diffeomorphism; a differentiable invertible function with a differentiable inverse. Namely, given a simple prior density $q_{z}(z)$ and a parameterized flow (diffeomorphism) $f_{\theta}(x)$, one can define the push-forward distribution as the map of samples from the prior distribution $z\sim q_{z}(z)$ via the inverse flow $x=f^{-1}_{\theta}(z)\sim q_{\theta}(x)$ with learnable parameters $\theta$. The density of the push-forward distribution can then be computed via the change-of-variables formula
@@
Autoregressive normalizing flows (Kingma et al., 2016; Papamakarios et al., 2017; Zhai et al., 2024) define a family of invertible maps with tractable Jacobian as a sequence of composed transformations $f_{\theta}=f_{\tau}\circ\ldots\circ f_{0}$, where each transformation $z_{t+1}=f_{t}(z_{t})$, $z_{0}=x$ is defined autoregressively. In the case of the TarFlow (Zhai et al., 2024), this is an autoregressive affine update defined over blocks of latent variable corresponding to image patches $z_{t}[i]\in\mathbb{R}^{D}$. That is, the $i$-th latent block is
@@
where we adopt slicing notation denoting the $i$-th block as $z[i]$ and blocks up to the $i$-th (exclusive) as $z[:i]$. Notably, the autoregressive structure allows for efficient evaluation of the Jacobian determinant due to its lower-triangular structure $\log|\partial f_{t}(z_{t})/\partial{z_{t}}|=-\sum_{i=1}^{N-1}\sum_{j=0}^{D-1}\alpha_{t}(z_{t}[:i])[i]_{j}$. Furthermore, the autoregressive affine updates are invertible with inverse $z_{t}=f^{-1}_{t}(z_{t+1})$ given by
$z_{t}[i]=\begin{cases}z_{t+1}[i]\,,&i=0\,,\\ z_{t+1}[i]\odot\exp(\alpha_{t}(z_{t}[:i])[i])+\mu_{t}(z_{t}[:i])[i]\,,&i\in[1,N-1]\,.\end{cases}$ (3)
However, clearly, such transformations leave the leading dimension $z_{t}[0]$ untouched, hence must be interleaved with permutations $\pi_{t}$ over the dimensions $f_{\theta}=\pi_{\tau}\circ f_{\tau}\circ\ldots\circ\pi_{0}\circ f_{0}$. For example, Zhai et al. (2024) use simple inversions of the latent block sequence for all $\pi_{t}$ across the entire model.
@@ 2.2 Boltzmann generators
Despite normalizing flows allowing various forms of training supervision, such as variational inference for unnormalized densities (Rezende and Mohamed, 2015), or maximum likelihood from empirical distributions (Kingma and Dhariwal, 2018), errors present in the parameterized distribution prevents accurate evaluation within scientific applications requiring high precision, e.g. free energy estimation. Boltzmann generators address specifically this challenge by performing self-normalized importance sampling (SNIS) at inference time. Namely, to evaluate the expectation of a statistic $\varphi(x)$ w.r.t. the target Boltzmann density $p(x)$ one can use the following consistent Monte Carlo estimator
@@
where $q_{\theta}(x)$ is the density of the learned normalizing flow. The SNIS estimator converges, for $n\to\infty$, to the true value $\mathbb{E}_{p(x)}\varphi(x)$. Furthermore, Tan et al. (2025) extended the Boltzmann generator framework to more general Monte Carlo algorithms, in particular a continuous-time formulation of Sequential Monte Carlo (Jarzynski, 1997; Albergo and Vanden-Eijnden, 2025).
Transferable Boltzmann generator (TBG) (Klein and Noe, 2024) made a first attempt of learning a sampler that generalizes across target densities corresponding to different peptide systems. TBG parametrizes the proposal distribution as a continuous normalizing flow (CNF) (Chen et al., 2018b), where the vector field is defined by an equivariant graph neural network (Satorras et al., 2021; Klein et al., 2023c). Crucial to the method is the system-dependent conditioning of $N$ atoms
$h[i]=[A_{i},R_{i},P_{i}]\,$, (5)
where atom type $A_{i}$, residue type $R_{i}$, and residue position $P_{i}$ are each encoded as one-hot vectors. Training on a set of MD trajectories for dipeptide systems (Klein et al., 2023b) with this system-conditional encoding enables TBG to generate a proposal for previously unseen dipeptides.
However, despite successful generalization across dipeptides, the TBG architecture introduces significant bottlenecks for inference and fine-tuning. Indeed, the learned CNF requires accurate integration of the vector field and computationally expensive evaluation of its divergence for evaluating the learned density model. For instance, the implementation of (Klein and Noe, 2024) requires 4 GPU-days to produce $3\times 10^{4}$ samples with their corresponding proposal likelihoods for a single dipeptide system. Furthermore, the expensive evaluation of density makes it infeasible to train or finetune TBG via the reverse KL-divergence or create a replay buffer of a substantial size.
@@ 3.1 Architecture of Prose
Prose builds on the TarFlow architecture (Zhai et al., 2024), which parametrizes a sequence of autoregressive affine transformations via blocks of transformer layers. The expressivity and favorable scalability of the transformer layers enables TarFlow to effectively model high dimensional data, whilst the affine autoregressive flow parameterization ensure fast and accurate energy evaluation. With minimal modifications TarFlow is capable of successfully modeling high-dimensional molecular data (Tan et al., 2025). Here we describe our design choices that make transferability possible.
@@
We extend TarFlow to support concurrent training on sequences of variable length. Whilst transformers natively support sequences of arbitrary length, special consideration is required within a normalizing flow such as TarFlow that is defined for fixed input and output dimensions. We therefore define appropriate masking to the affine sequence updates and log-determinant aggregation to prevent padding tokens influencing either computation, under arbitrary sequence permutations; further details of the tokenization and masking are provided in Appendix˜A. We additionally replace the fixed-length learnable position embedding with the more extrapolation-friendly sinusoidal embedding. This design enables Prose to efficiently train across a distribution of systems $s$ by maximizing the normalized log-likelihood
@@
where $d(s)$ is the size of the system $s$ (Klein and Noe, 2024). This extended architecture allows for parallel processing of data dimensions, enabling transferability and scalability across lengths.
@@
The standard TarFlow employs simple additive conditioning for class-conditional image generation. Whilst we find this to be sufficient to define a system-transferable normalizing flow, we follow large-scale atomistic transformer architectures in applying conditioning through adaptive layer normalization, adaptive scaling, and SwiGLU transition blocks (Abramson et al., 2024; Geffner et al., 2024). The system conditioning features are constructed from atom types $A$, residue types $R$, residue positions $P$, and sequence lengths $L$. Atom and residue types are embedded using lookup-table embedding layers, whilst sinusoidal embeddings are employed for the naturally ordered sequence position and sequence length. See Appendix˜A for further details.
@@
In the image setting, Zhai et al. (2024) employ only an identity and flip permutation to the sequence of image tokens. Similarly, when applying TarFlow to peptide systems Tan et al. (2025) employ only an identity and flip permutation on the ordering defined by the PDB format, in which the sequence is defined per-residue starting with backbone atoms followed by sidechain atoms. Whilst a simple identity and flip may be appropriate for the regular grid of image data, we argue this to be suboptimal for the diversity of geometric interactions present in molecular systems. This motivates our introduction of chemistry-aware sequence permutations, defined to promote effective peptide modeling. We define the backbone permutation, such that the backbone atoms $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=0}^{L-1}$ (with associated hydrogens) for all residues are located at the start of the sequence, and followed by the sidechains. By processing the coordinates of the backbone atoms at the start of the sequence, the model refines the global structure of the peptide as a contiguous sequence. Crucially, the sidechain positions are subsequently able to causally attend to the full backbone structure, hence enabling local updates to be influenced by global structure. We further employ a backbone-flip permutation to provide additional diversity to the autoregressive modeling.
Figure 2: All-atom block-wise autoregressive normalizing flow based on the TarFlow (Zhai et al., 2024). Peptides are encoded via atom types $A$, residue types $R$, residue position $P$, and sequence length $L$. Atom positions in 3D Cartesian coordinates define the system state. The embedding of the peptide is applied as conditioning to the coordinates such that Prose achieves transferability between systems. Within each block the sequence $z_{t}$ is permuted and passed to a transformer, defining an autoregressive affine update. In the backbone permutation the backbone $N_{i},C_{\alpha,i},C_{i},O_{i}]_{i=0}^{L-1}$ of all residues (with associated hydrogens) is updated before any sidechains, providing additional diversity to the causal attention for global structure modeling.
@@ 3.2 Inference and fine-tuning of Prose
In practice, we compare the true density $p(x)$ with our generated distribution $\tilde{p}(x)$ instead of measuring statistics $\varphi(x)$. For completeness, we describe all the considered estimators in Appendix˜D.
@@ 4.2 Experimental configuration
We train the first Boltzmann generators transferable across peptide sequence length. We train the Prose architecture defined in Section˜3.1, an unmodified TarFlow (Zhai et al., 2024) as in SBG (Tan et al., 2025), and the equivariant CNF of Klein and Noe (2024), with the improved training recipe of Tan et al. (2025), denoted as ECNF++. All models are trained for $5\times 10^{5}$ iterations with batch size 512. Both Prose and TarFlow are suitably scalable to long sequences and are trained on the full dataset detailed in Section˜4.1. However, generating 8 residue sequences with likelihoods for ECNF++ was found to be prohibitively expensive, hence the training data was limited to sequences up to and including length 4. Comprehensive training details are provided in Appendix˜C.
@@ 4.3 Scale transferability of Prose
To establish the performance of Prose as a sampler proposal distribution, we first evaluate the trained flows in the Boltzmann generator setting. Here we generate a set of proposal particles $\{x_{i}\}_{i=1}^{N}$, evaluate model likelihoods $q_{\theta}(x_{i})$ and reweight using SNIS as in Eq.˜7. In addition to the trained models we benchmark against the following pretrained baselines; (i) the TBG model trained by Klein and Noe (2024), denoted as ECNF (ii) TimeWarp (Klein et al., 2023a) (iii) BioEmu (Lewis et al., 2024) (iv) Unisim (Yu et al., 2025). For all methods we permit a budget of $10^{4}$ energy evaluations. For the Boltzmann generator methods (ECNF, ECNF++, TarFlow, Prose) this corresponds to $10^{4}$ SNIS particles; for further information on the budget allocation of non-BG methods (TimeWarp, BioEmu, UniSim) see Appendix˜E. We additional provide results for the unweighted proposal distributions in Appendix˜F. We note the TimeWarp dataset to lack any sequences containing Proline at the N-Terminal, hence neither TimeWarp nor ECNF were evaluated on such sequences.